OpenStudy (anonymous):

find the greatest common factor of the following monomials: 46a^2 28a 20a^6

OpenStudy (cwrw238):

find prime factors: 46 = 2 * 23 28 = 2 * 2 * 7 20 = 2 * 2 * 5 the GCF = number common to all 3 sets of factors = 2 GCF of terms in a = a GCF of three terms = 2a

OpenStudy (anonymous):

thanx :)

OpenStudy (anonymous):

thanx :)

OpenStudy (cwrw238):

yw